OpenStudy (zacharydoc):

How many solutions does the following equation have? |4x + 12| = 0 No solution One solution Two solutions Infinitely many solutions

OpenStudy (zacharydoc):

help

OpenStudy (shansha1):

one solution

OpenStudy (zacharydoc):

how did u get that

OpenStudy (zacharydoc):

\[\frac{ 3x-15 }{ 2}=4x\]

OpenStudy (shansha1):

thats the wrong one thats for someone else sorry

OpenStudy (zacharydoc):

what about this problem

OpenStudy (zacharydoc):

its above your last comment

OpenStudy (shansha1):

ight

OpenStudy (zacharydoc):

\[\frac{ 3x-15 }{ 2}=4x\]

OpenStudy (shansha1):



OpenStudy (shansha1):

x=2

OpenStudy (zacharydoc):

thanks

OpenStudy (shansha1):

ur welcome